Hosts A and B are communicating over a TCP connection, and Host B has already received from A all bytes up to and including byte 1029. Suppose Host A then sends three segments to Host B back-to-back. The first, second and third segments contain 1200, 150 and 80 bytes of data respectively. In the first segment, the sequence number is 1060, the source port number is 8080, and the destination port number is 80. Host B sends an acknowledgment whenever it receives a segment from Host A. Answer the following questions:

A) What is the sequence number in the second segment sent by Host A?
B) What is the sequence number in the header of the third segment sent by Host A
C) What is the destination port number in the header of the third segment sent by Host A?
D) If the first segment arrives before the second and third segments, what is contained in the acknowledgment number field of the segment that B sends to A? (Provide answer in decimal value)
E) If the third segment arrives before the first and second segments, what is in the ACK number field of the segment that B sends to A in response to this (third) segment?
Answer #1

A) sequence number sent by host A = 1060+1200=2260

as 1060 is the last sequence number and 1200 bytes of data is send .

B) sequece number in header of 3rd segment = 2260+150=2410

C)destination port = 80

D) acknowledgement number = 2260

this shows it is still waiting for byte number 2260 i.e. second frame

E) acknowledgement number = 1060

this indicates it is still waiting for the first segment